# Disentangling Task Interference within Neurons: Model Merging in Alignment with Neuronal Mechanisms

### v1

@@ 3.1 Preliminaries
3.2 Neuronal Task Vector
Zooming in to the neuronal level of $\tau_{t}$, we define $\mathbf{w}^{k}_{t}\in\mathbb{R}^{d}$ to represent the weight vector of $k^{th}$ neuron in model $\theta_{t}$. The task-specific adaptation of this neuron, relative to the pre-trained model, is given by the neuronal task vectors:
$\mathbf{\tau}^{k}_{t}=\mathbf{w}^{k}_{t}-\mathbf{w}^{k}_{0},$ (1)
where $\mathbf{w}^{k}_{0}$ corresponds to the weight of the $k^{th}$ neuron in the pre-trained model $\theta_{0}$. Figure 2 illustrates our proposed neuronal task vector.
3.3 Neuronal Subspace Decomposition
To examine how task-specific modifications impact individual neurons, we decompose the neuronal task vectors into two complementary neuronal subspaces, visualized in Figure 2. Mathematically, this decomposition is formulated as:
$\tau^{k}_{t}=\tau^{k}_{\parallel,t}+\tau^{k}_{\perp,t},$ (2)
where $\tau^{k}_{\parallel,t}=\mathbf{P}\tau^{k}_{t}$ projects the neuronal task vectors onto the pre-trained model’s weight space, Parallel Subspace ($\mathcal{P}$), capturing neuron sensitivity. Here, $\mathbf{P}$ is the projection matrix onto the span of $\mathbf{W}^{k}_{0}$. $\tau^{k}_{\perp,t}=(\mathbf{I}-\mathbf{P})\tau^{k}_{t}$ captures the complementary orthogonal modifications, in the Orthogonal Subspace ($\mathcal{O}$), reflecting task-specific adaptability. The role of each complementary subspace:
The impact of weight editing in these complementary subspace on task adaption is discussed in Section 6.3.
3.4 NeuroMerging
Subsequently, we reconstruct the task vector $\overline{\tau}$ with the all the merged neuronal task vectors $\overline{\tau}_{k}$. Finally, we obtained the final merged model $\overline{\theta}=\theta_{0}+\overline{\tau}_{rescaled}$.
4 Experimental Setup
Baseline Methods. Our baselines comprise two main categories: (1) non-model merging approaches, which include individually fine-tuned models and a multitask model trained jointly on the combined dataset serving as our theoretical upper bound, and (2) various advanced model merging techniques, including Simple Averaging (Wortsman et al., 2022), Fisher Merging (Matena and Raffel, 2022), RegMean (Jin et al., 2023), Task Arithmetic (Ilharco et al., 2023), TIES-Merging-Merging (Yadav et al., 2024), PCB-Merging (Du et al., 2024), and our proposed NeuroMerging method. We report average accuracy across all tasks’ test sets as our primary evaluation metric.
Notably, task vector-based approaches such as Task Arithmetic, Ties-Merging, and PCB-Merging, along with our proposed NeuroMerging, are substantially more lightweight and efficient. These approaches are training-free and do not rely on a validation set, making them highly practical for real-world applications. To further evaluate this advantage, we conducted additional experiments comparing task vector-based methods in scenarios where validation sets were unavailable.
Task($\rightarrow$) Test Set Performance
Zeroshot - 53.1 58.2 54.2 54.1 54.3 70.9 49.2 63.9
Finetuned - 88.0 94.4 97.1 85.3 91.0 95.7 71.6 80.6
Multitask - 88.1 94.2 98.5 89.3 92.0 95.4 73.5 73.6
Averaging[ICML22] 51.6 59.2 26.3 69.6 53.8 67.3 49.1 36.1
Task Arithmetic[ICLR23] 59.7 60.9 31.7 57.8 73.0 73.5 55.7 65.3
TIES-Merging[NeurIPS23] 76.7 80.8 92.4 77.7 81.9 78.4 61.9 63.9
PCB-Merging[NeurIPS24] 76.9 82.9 93.2 79.0 84.4 75.6 63.5 59.7
NeuroMerging (Ours) 77.6 81.1 94.3 81.6 84.7 81.2 56.7 83.9
Fisher Merging[NeurIPS22] 68.7 68.4 83.0 65.5 62.4 94.1 58.2 49.2
RegMean[ICLR23] 79.8 83.9 97.2 73.2 82.6 94.1 63.2 64.4
Task Arithmetic[ICLR23] 74.6 72.7 91.3 76.4 85.6 74.4 61.0 61.1
TIES-Merging[NeurIPS23] 79.5 82.6 94.9 72.8 87.4 85.2 66.6 66.7
PCB-Merging[NeurIPS24] 81.0 87.0 95.2 76.4 88.1 88.4 64.3 68.1
NeuroMerging (Ours) 82.2 86.4 94.3 75.9 87.9 91.2 65.9 73.6
5 Results
5.1 Merging NLP Models
Following the experimental settings from Yadav et al. (2024), we use the T5-base and T5-large models (Raffel et al., 2020), which are encoder-decoder transformers (Vaswani et al., 2017) pretrained via masked language modeling on a large text corpus, and fine-tune them independently on seven tasks: Khot et al.’s (2019) QASC, Yang et al.’s (2015) WikiQA, and Tafjord et al.’s (2019) QuaRTz for Question Answering; Zhang et al.’s (2019) PAWS for Paraphrase Identification; Sharma et al.’s (2018) Story Cloze for Sentence Completion; and Sakaguchi et al.’s (2021) Winogrande together with Levesque et al.’s (2012) WSC for Coreference Resolution. Table 2 and 7 demonstrated that our approach has superior performance over state-of-the-art methods, achieving improvements of $0.6\%$ and $1.2\%$ for T5-base and T5-large, respectively. Moreover, NeuroMerging without validation showed even more substantial gains, surpassing previous methods by $1.4\%$ for T5-base and $0.7\%$ for T5-large. For comprehensive results across all tasks and model variants, see Appendix B.2.
Figure 3: In-domain v.s. Out-domain performance.
5.2 Out-of-Domain Generalization
Building upon the experimental setup of Yadav et al. (2024), we also investigated how merged models between tasks enhance generalization in different domains. Following the approach used in prior NLP models, we merged models on seven in-domain datasets and evaluated their performance on six held-out datasets from the T0 mixture (Sanh et al., 2022) to assess cross-task generalization. These datasets encompass diverse tasks, covering three Question Answering datasets: Huang et al.’s (2019) Cosmos QA, Sap et al.’s (2019) Social IQA, and Rogers et al.’s (2020) QuAIL; one Word Sense Disambiguation dataset: Pilehvar and Camacho-Collados’s (2019) WiC; and two Sentence Completion datasets: Gordon et al.’s (2012) COPA and Zellers et al.’s (2019) H-SWAG. As shown in Figure 3, NeuroMerging outperformed the strongest baseline by $0.7\%$ and $0.5\%$ for T5-base and T5-large models, respectively, showcasing enhanced out-of-domain generalization capabilities. For more comprehensive results, please refer to Appendix B.2 Table 10 and 11.
5.3 Merging LLMs
We follow the experimental setup of Du et al. (2024) and extend our NeuroMerging to larger LLMs. Specifically, we merged three domain-specialized Llama-2-7b models (Touvron et al., 2023), and each is fine-tuned for distinct capabilities: Chinese language understanding [^1] https://huggingface.co/LinkSoul/Chinese-Llama-2-7b, mathematical reasoning [^2] https://huggingface.co/meta-math/MetaMath-7B-V1.0 (Yu et al., 2024), and code generation [^3] https://huggingface.co/qualis2006/llama-2-7b-int4-python-code-18k (Roziere et al., 2023). To rigorously evaluate the performance of each specialized model, we employed established benchmarks tailored to their respective domains: Li et al.’s (2024) CMMLU for assessing Chinese language proficiency, Cobbe et al.’s (2021) GSM8K for mathematical capabilities, and Chen et al.’s (2021) HumanEval for code generation competency. As demonstrated in Table 3, our NeuroMerging approach exhibited substantial performance improvements, surpassing the strongest baseline by $0.9\%$. Notably, our method even outperformed the PCB-Merging utilizing Evolution Strategies (ES) optimization algorithm by $0.7\%$, underscoring the effectiveness of our proposed methodology.
Table 3: Performance comparison on LLaMA2.
Model CMMLU GSM8K Human-Eval Average
Chinese 38.6 2.3 13.4 18.1
Math 31.2 65.6 0.0 32.3
Code 33.3 0.0 17.1 16.8
Averaging[ICML22] 35.6 48.5 6.7 30.3
Task Arithmetic[ICLR23] 35.4 46.1 9.8 30.4
TIES-Merging[NeurIPS23] 36.5 53.4 12.8 34.3
PCB-Merging[NeurIPS24] 36.4 52.3 16.5 35.1
PCB-Merging+ES [NeurIPS24] 36.4 53.1 16.5 35.3
NeuroMerging (Ours) 36.1 57.2 14.6 36.0
Table 4: Performance comparison on ViT.
Method Validation ViT-B/32 Avg. ViT-L/14 Avg.
Individual - 90.5 94.2
Multi-task - 88.9 93.5
Averaging[ICML22] 65.8 79.6
Task Arithmetic[ICLR23] 60.4 83.3
TIES-Merging[NeurIPS23] 72.4 86.0
PCB-Merging[NeurIPS24] 75.9 86.9
NeuroMerging (Ours) 76.4 87.9
Fisher Merging[NeurIPS22] 68.3 82.2
RegMean[ICLR23] 71.8 83.7
Task Arithmetic[ICLR23] 70.1 84.5
TIES-Merging[NeurIPS23] 73.6 86.0
PCB-Merging[NeurIPS24] 76.3 87.5
NeuroMerging (Ours) 76.5 88.3
5.4 Merging Vision Models
6 Additional Results and Analysis
6.1 Merging without Validation Sets
When validation data is unavailable, we examine the parameters $\lambda_{1}$ and $\lambda_{2}$ selected according to Section 3.4, where $\lambda_{1}$ is set to zero, as the corresponding subspace is observed to have little impact in Section 6.4. The value of $\lambda_{2}$ is computed based on the L1-norm of masked and unmasked task vectors. Figure 7 and Table 2 and Appendix Tables 2,  8, and 9 present the evaluation of NeuroMerging on NLP and CV tasks across various model sizes, comparing it with existing methods. NeuroMerging achieves the highest average accuracy across all tasks. This demonstrates that our proposed method, with a simple rescaling, outperforms existing methods on average. Specifically, it achieved a 1.4% and 0.7% improvement over the strongest baseline for T5-Base and T5-Large, respectively. For vision models, it outperforms the strongest baseline by 0.5% and 1.0% for ViT-B/32 and ViT-L/14, respectively.
Table 5: Comparison of $\psi(\cdot)$ on Avg. Accuracy.
@@ 6.1 Merging without Validation Sets
6.2 Ablation Study on Merging Functions
6.3 Role of Neuronal Subspaces
Table 6: In/Out-domain Performance of T5-large.
T5-large In-domain Out-domain Total Average
Fine-tuned 88.0 53.8 58.7
Keep Orthogonal 88.0 53.9 58.8
Keep Parallel 52.7 51.8 51.9
6.4 Robustness of Hyperparameters
Figure 4: Impacts of $\lambda_{1}$ and $\lambda_{2}$ on T5-Large.
7 Conclusion
In this paper, we revisited model merging from the core principle neuron and its connectivity that underpin the learning process of recent deep neural networks and LLMs. Specifically, we presented the first exploration into the impact of neuronal alignment in the model merging process, by decomposing of task-specific representations into two complementary neuronal subspaces and examining their roles. Based on these insights, we proposed NeuroMerging, a novel framework designed to reduce task interference within neurons. Our empirical evaluations demonstrate that NeuroMerging achieves superior performance compared to existing approaches on multi-task benchmarks across both vision and natural language processing tasks.
Future work could extend NeuroMerging to larger models or multimodal architectures with more tasks. While our study introduces a neuronal mechanistic perspective on task interference in model merging, it captures only one aspect of neuronal mechanisms, as it does not yet explore interactions between synchronized neuron groups or neuronal dynamics. Further research is needed to investigate these factors and deepen our understanding of neuronal mechanisms in model merging or even multi-task learning.
8 Limitations
While our work provides the first neuronal mechanistic perspective on task interference when merging language models, (1) it remains a partial view of neuronal mechanisms as it does not yet explore interactions between synchronized neuron groups or neuronal dynamics during merging and inference, which require further investigation. Moreover, this work shares similar limitations with current SOTA model merging methods, including (2) the effectiveness of task arithmetic in model merging relies on selecting fine-tuned checkpoints that are beneficial for specific domains, ensuring they originate from the same pretrained model, and addressing hyperparameter sensitivity; and (3) More effort is needed to develop a mathematical understanding of why and when task arithmetic in model merging works well, despite its simplicity and efficiency.
Appendix A Statistical Analysis of Parameters
Appendix B Additional Results
B.1 Mask Ratio vs. Performance
@@ Appendix B Additional Results
B.2 Comprehensive Task-Level Results
We provide all task-level results in T5-Base, T5-Large (Raffel et al., 2020), LLaMA2 (Touvron et al., 2023), ViT-B/32, and ViT-L/14 (Dosovitskiy et al., 2021), respectively. The task-level results of the in-domain experiments for all models can be found in Tables 789. The task-level results of the out-domain experiments for T5-Base and T5-Large can be found in Tables 10 and 11. Lastly, Table 12 shows the task-level results from Section 6.3 when only one of the neuronal subspace is retained.
@@ B.2 Comprehensive Task-Level Results
Table 12: Test set performance comparison of T5-Large models under different keeping strategies (naive finetuned, keep orthogonal, and keep parallel) across seven NLP tasks. Please refer to Section 1 and 6.3 for experimental details.
@@
Appendix C Implementation Details
Appendix D Dataset Details

### v2

@@ Appendix A Additional Details
A.1 Checkpoints for Llama
We merged three domain-specialized Llama-2-7b models (Touvron et al., 2023), and each is fine-tuned for distinct capabilities: Chinese language understanding [^1] https://huggingface.co/LinkSoul/Chinese-Llama-2-7b, mathematical reasoning [^2] https://huggingface.co/meta-math/MetaMath-7B-V1.0 (Yu et al., 2024b), and code generation [^3] https://huggingface.co/qualis2006/llama-2-7b-int4-python-code-18k (Roziere et al., 2023).
Algorithm 1 is the pseudo-code for NeuroMerging.
A.2 Masking
As entries of the task vector with larger magnitudes are more relevant to task-specific adaptation, only the top $r\%$ of $\mathbf{\tau}_{t}$ with the largest magnitudes are kept, while the others are set to zero Yadav et al. (2023). The masked task vector is defined as $\tau^{\text{masked}}_{t}=m_{t}\circ\tau_{t}$, where $m_{t}$ is the mask that keeps the top $r\%$ of the elements of each task vector. In the main manuscript, we use $\tau_{t}$ to represent the masked task vector for readability.
Appendix B Why Merging at the Neuronal Level?
Moreover, we demonstrated in the main manuscript that merging in the orthogonal subspace ($\mathcal{O}$) is more effective. To explain why, we mathematically show that $\bar{x}_{\perp}$ does not influence the activation of the pre-trained neuron. For comparison, consider the activation of the pre-trained neuron:
$y=\phi(\mathbf{w}^{k}_{0}\cdot\bar{x})=\phi(\mathbf{w}^{k}_{0}\cdot(\bar{x}_{\parallel}+\bar{x}_{\perp}))$ (8)
Since $\bar{x}_{\perp}$ belongs to the complementary orthogonal null space of $\mathbf{w}^{k}_{0}$, we have:
$\mathbf{w}^{k}_{0}\cdot\bar{x}_{\perp}=0$ (9)
Due to vector product of orthogonal subspaces, we simplify the activation as:
$y=\phi(\mathbf{w}^{k}_{0}\cdot\bar{x}_{\parallel})$ (10)
Hence, the orthogonal subspace ($\mathcal{O}$) encapsulates more task-specific adaptations, making it a more effective space for merging. In Table 6, we summarized the differences between our work with existing methods.
Method Scale Granularity Level
Fisher Merging[NeurIPS22] Fisher Matrix Parameter
RegMean[ICLR23] Inner Product Matrix Parameter
Task Arithmetic[ICLR23] Uniformed Task
Ties-Merging[NeurIPS23] Uniformed Parameter
DARE[ICML24] $1/(1-p)$ Parameter
LoraHub[COLM24] Evolver Searched Task
AdaMerging[ICLR24] Unsupervised Optimized Layer
PCB-Merging[NeurIPS24] Balancing Matrix Parameter
NeuroMerging (Ours) L1-Norm Neuron
Table 6: Model merging with different scales and granularity levels.
Appendix C Additional Results
C.1 Ablation Study on Merging Functions
@@
Table 7: Comparison of $\psi(\cdot)$ on Avg. Accuracy.
C.2 Mask Ratio vs. Performance
@@
Task($\rightarrow$) T5-Base T5-Large
Method($\downarrow$) Acc. (%) Orth. ($\mathcal{O}$) Par. ($\mathcal{P}$) Acc. (%) Orth. ($\mathcal{O}$) Par. ($\mathcal{P}$)
TIES-Merging[NeurIPS2023] 76.4 76.2 55.5 79.5 79.9 52.5
PCB-Merging[NeurIPS2024] 76.9 77.1 54.7 81.0 81.8 52.0
NeuroMerging (Ours) 77.5 77.5 54.6 82.2 82.2 52.2
Table 8: NLP Task Accuracy under Orthogonal and Parallel Projections across T5-Large and T5-Base Models
Task($\rightarrow$) ViT-B/32 ViT-L/14
Method($\downarrow$) Acc. (%) Orth. ($\mathcal{O}$) Par. ($\mathcal{P}$) Acc. (%) Orth. ($\mathcal{O}$) Par. ($\mathcal{P}$)
AdaMerging[ICLR2024] 71.1 71.8 48.9 84.9 85.0 65.1
AdaMerging++[ICLR2024] 73.7 74.5 49.2 87.3 87.5 65.6
PCB-Merging[NeurIPS2024] 76.3 76.4 48.7 87.5 87.8 65.0
NeuroMerging (Ours) 76.5 76.6 48.7 88.3 88.3 65.1
Table 9: Vision Task Accuracy under Orthogonal and Parallel Projections across ViT-B/32 and ViT-L/14 Models
Task($\rightarrow$) T5-Base T5-Large
Method($\downarrow$) Time (s,$\downarrow$) Mem (GB,$\downarrow$) Acc. (%$\uparrow$) Time (s,$\downarrow$) Mem (GB,$\downarrow$) Acc. (%$\uparrow$)
TIES-Merging[NeurIPS2023] 17 61.4 76.4 50 190.1 79.5
PCB-Merging[NeurIPS2024] 98 105.5 76.9 323 312.1 81.0
NeuroMerging (Ours) 67 18.7 77.5 161 69.1 82.2
Table 10: Comparison on Computational Efficiency and Accuracy for NLP Tasks
Task($\rightarrow$) ViT-B/32 ViT-L/14
Method($\downarrow$) Time (s,$\downarrow$) Mem (GB,$\downarrow$) Acc. (%$\uparrow$) Time (s,$\downarrow$) Mem (GB,$\downarrow$) Acc. (%$\uparrow$)
TIES-Merging[NeurIPS2023] 12 28.4 73.6 36 84.7 86.0
PCB-Merging[NeurIPS2024] 51 45.4 76.3 149 137.1 87.5
AdaMerging[ICLR2024] 6793 80.1 71.1 18871 265.4 84.9
AdaMerging++[ICLR2024] 6384 95.7 73.1 18184 335.7 87.3
NeuroMerging (Ours) 67 8.8 76.5 141 21.7 88.3
Table 11: Comparison on Computational Efficiency and Accuracy for Vision Tasks
C.3 Comprehensive Task-Level Results
We provide all task-level results in T5-Base, T5-Large (Raffel et al., 2020), LLaMA2 (Touvron et al., 2023), ViT-B/32, and ViT-L/14 (Dosovitskiy et al., 2021), respectively. The task-level results of the in-domain experiments for all models can be found in Tables 121314. The task-level results of the out-domain experiments for T5-Base and T5-Large can be found in Tables 15 and 16. Lastly, Table 17 shows the task-level results from Section 6.2 when only one of the neuronal subspace is retained.
C.4 Existing Model Merging Methods in Neuronal Subspaces
To empirically validate the neuronal subspaces, we extended our experiments to various existing model merging methods. As shown in Tables 8 and 9, experimental results demonstrate that applying these merging methods within our proposed orthogonal neuronal subspaces consistently enhances their performance. Notably, this improvement occurs despite these existing methods not explicitly incorporating neuronal mechanisms into their designs. This observation further underscores the significance and motivation of our proposed approach.
C.5 Scalability Evaluation
We further evaluated the scalability of NeuroMerging from the perspective of computational efficiency. As summarized in table 10 and 11 (runtime in seconds, memory usage in GB; lower is better), NeuroMerging consistently achieves superior efficiency compared to existing methods across representative NLP and CV benchmarks. Specifically, it demonstrates comparable or reduced memory usage and improved accuracy relative to TIES-Merging Yadav et al. (2023) ($+2.7\%$ on T5-Large, $+2.3\%$ on ViT-L/14, $+1.7\%$ on LLaMA2), significantly lower runtime and memory overhead compared to PCB-Merging Du et al. (2024), and vastly reduced computational requirements relative to AdaMerging Yang et al. (2024b), which incurs up to $100$ times longer runtime and $10$ times higher memory usage on ViT-L/14 due to additional training steps. In general, unlike prior approaches Yadav et al. (2023); Du et al. (2024); Yang et al. (2024b); Jung et al. (2024); Yang et al. (2024a); Huang et al. (2024b); Lu et al. (2024); Matena and Raffel (2022), which either require loading all parameters simultaneously or introduce costly auxiliary components, NeuroMerging leverages neuronal task vectors aligned with neuronal mechanisms, ensuring scalable, lightweight, and efficient merging.
Appendix D Implementation Details
Appendix E Dataset Details
@@
Table 17: Test set performance comparison of T5-Large models under different keeping strategies (naive finetuned, keep orthogonal, and keep parallel) across seven NLP tasks. Please refer to Section 1 and 6.2 for experimental details.